### v1

# MLVTG: Mamba-Based Feature Alignment and LLM-Driven Purification for Multi-Modal Video Temporal Grounding

Abstract
Video Temporal Grounding (VTG), which aims to localize video clips corresponding to natural language queries, is a fundamental yet challenging task in video understanding. Existing Transformer-based methods often suffer from redundant attention and suboptimal multi-modal alignment. To address these limitations, we propose MLVTG, a novel framework that integrates two key modules: MambaAligner and LLMRefiner. MambaAligner uses stacked Vision Mamba blocks as a backbone instead of Transformers to model temporal dependencies and extract robust video representations for multi-modal alignment. LLMRefiner leverages the specific frozen layer of a pre-trained Large Language Model (LLM) to implicitly transfer semantic priors, enhancing multi-modal alignment without fine-tuning. This dual alignment strategy, temporal modeling via structured state-space dynamics and semantic purification via textual priors, enables more precise localization. Extensive experiments on QVHighlights, Charades-STA, and TVSum demonstrate that MLVTG achieves state-of-the-art performance and significantly outperforms existing baselines.
Keywords:
Video Temporal Grounding Mamba Large Language Model
Figure 1: MLVTG projects features of video-text pairs into a shared semantic space, narrows the semantic gap, and better aligns visual-text features. Its dual-branch architecture handles temporal localization and highlight detection separately, allowing task-specific optimization.
1 Introduction
In video understanding research, untrimmed videos dominate digital content yet contain sparse valuable clips, creating critical demand for Video Temporal Grounding (VTG) - the task of aligning queries with precise temporal localization in untrimmed videos. Researchers consider that its fundamental capability covers two key tasks: Temporal Localization (TL) and Highlight Detection (HD). However, effectively executing these tasks presents significant challenges. Current limitations stem from inherent modality discrepancies in multi-modal alignment, where heterogeneous representations of visual dynamics and linguistic structures create matching ambiguities that degrade system performance [23]. Addressing these challenges could revolutionize human-AI video interaction, enabling real-time applications.
Early methods rely on 3D CNNs [9, 14, 7, 19], but limit receptive fields resulted in insufficient video representation capability, leading to coarse multi-modal alignment. Transformer-based models [4, 18, 26, 23], leveraging the attention mechanism, have advanced VTG by effectively capturing global contextual dependencies. However, Transformers often suffer from redundant attention across frames [42, 3, 1], weakening temporal discrimination. Recently, state space models(SSMs) based Mamba [11] have demonstrated strong sequence modeling capacity in vision tasks [25, 46] by dynamically filtering redundant signals. Yet, directly applying Mamba to VTG lacks temporal expressiveness. To address this, we adopt the bidirectional scanning strategy from [20, 46], stacking multiple Vision Mamba blocks to extract robust video representations for multi-modal alignment.
To ensure that video temporal modeling possesses stronger semantic expressiveness and accuracy, we further introduce large language mode (LLM) to enhance multi-modal alignment. It exhibits strong multi-modal reasoning abilities [34, 6, 41], effectively suppressing noise, purifying semantics and enhancing multi-modal alignment in video tasks [45, 5]. To leverage such capabilities, we utilize the specific frozen layer of pre-trained LLMs. With frozen pre-trained parameters, LLMs can transfer textual priors to visual domains, filtering redundant video content. According to the Platonic Representation Hypothesis [17], neural networks converge to a shared semantic space across modalities, allowing LLMs to associate abstract textual concepts (e.g., “running”) with temporal visual patterns (e.g., “leg movement”), thereby improving alignment and robustness.
Overall, we propose Mamba-based feature alignment and LLM-driven purification for multi-modal Video Temporal Grounding termed MLVTG, a novel framework that combines MambaAligner and mamba-based LLMRefiner. MLVTG features a dual-branch alignment strategy: structured SSM enables adaptive feature selection, while a frozen LLM layer performs semantic purification. By replacing the Transformer with vision mamba blocks, our model reduces redundant attention, enhances temporal modeling, and aligns multi-modality. The SSM branch captures salient video context via selective memory, and the LLMRefiner projects high-level semantics without fine-tuning. MLVTG is the first to jointly leverage Mamba and LLM for VTG, achieving SOTA results on three benchmarks. Our key contributions are:
  • We propose MLVTG, a novel dual-branch architecture that handles temporal localization and highlight detection separately, allowing task-specific optimization.
  • We propose the MambaAligner, a stack of vision mamba blocks that replaces Transformer to better capture video representations and enhance multi-modal alignment.
  • We introduce the mamba-based LLMRefiner, which uses the specific frozen LLM layer to purify semantics through implicit multi-modal interactions without fine-tuning.
  • Experiments on QVHighlights, Charades-STA, and TVSum show that our method MLVTG outperforms baseline and achieves state-of-the-art results.
2 Related Work
2.1 Video Temporal Grounding
Video Temporal Grounding (VTG), a key task linking visual content with semantics, has evolved into two sub-tasks: Temporal Localization (TL) and Highlight Detection (HD). TL focuses on grounding natural language queries to video clips, with early proposal-based methods [9, 14] suffering from low coverage, while proposal-free models [19, 44] improve efficiency via direct boundary regression. HD, in contrast, identifies salient video moments, progressing from ranking-based methods [13] to DETR-based detectors [26] with improved accuracy. Though traditionally studied separately, TL and HD share temporal reasoning goals. M-DETR [18] unified them using QVHighlights, inspiring subsequent work on audio integration [24], negative sample learning, and multi-task setups [36]. However, existing methods still struggle with fine-grained multi-modal alignment. To this end, we introduce MLVTG, which jointly optimizes low-level temporal structure by MambaAligner and high-level semantics by LLMRefiner, achieving improvements in both TL and HD tasks.
2.2 Mamba for Sequence Tasks
State Space Models (SSMs) represent long-range sequences with linear complexity [11, 12], offering an efficient alternative to Transformers by leveraging continuous state representations. However, traditional SSMs lack the ability to dynamically focus on task-relevant features across varying contexts. To address this, Mamba [10] introduces gated selective SSMs, improving both noise suppression and salient temporal pattern extraction.
Originally surpassing Transformers in language modeling and time-series forecasting [10, 8, 22], Mamba has recently been extended to vision task  [46, 29] and video understanding [20, 27, 25]. However, existing Mamba-based models primarily focus on unimodal tasks and overlook fine-grained video-language alignment. To overcome this, we incorporate a bidirectional SSM mechanism inspired by VideoMamba, and stack multiple Vision Mamba blocks to capture robust video representations. This enables precise localization for VTG tasks by enhancing temporal modeling and multi-modal alignment.
2.3 Large Language Model for Video Temporal Grounding
Applying Large Language Models (LLMs) to VTG has become a focus of recent research, with several studies exploring different integration strategies. For example, VTimeLLM [16] employs LLM as the VTG decoder to directly output start and end timestamps. However, these existing methods face significant challenges, including limited capabilities in fine-grained temporal reasoning and precise timestamp localization inherent to current LLM architectures. Additionally, the requirement for extensive training exacerbates computational overhead. Motivated by [28], we propose the mamba-based LLMRefiner, a module that efficiently exploits semantic priors embedded within the specific frozen LLM layer. This strategy achieves enhanced semantic alignment between linguistic concepts and corresponding video clips, thereby significantly improving VTG performance.
3 Methodology
3.1 Problem Definition
VTG aims to align natural language descriptions with corresponding moments in untrimmed videos. It covers two major tasks: Temporal Localization and Highlight Detection, both requiring effective multi-modal reasoning.
3.1.1 Temporal Localization.
Given a video $\{f_{t}\}_{t=1}^{T}$ with frames $f_{t}\in\mathbb{R}^{H\times W\times 3}$, and a language query $\{q_{i}\}_{i=1}^{L_{q}}$, the goal is to predict a temporal clip $\tau=[\tau_{s},\tau_{e}]$ that best matches the query. The core challenge lies in fine-grained multi-modal alignment. A model $F_{\theta}(\{f_{t}\}_{t=1}^{T},\{q_{i}\}_{i=1}^{L_{q}})$ is trained to output $\hat{\tau}$ such that $\text{IoU}(\hat{\tau},\tau_{\text{gt}})\geq\alpha$, where $\alpha=0.5$ is the standard threshold.
3.1.2 Highlight Detection.
Given a video $\{f_{t}\}_{t=1}^{T}$, the objective is to assign a saliency score $s_{i}$ to each temporal clip $v_{i}$ and select the top-$K$ clips with the highest scores as highlights. Formally, the output is the set $M=\{f_{i}\mid s_{i}\in\text{top-}K\}$. In scenarios lacking explicit annotations, weak supervision signals such as video titles or domain labels can guide model training. [30, 32]
3.2 Overall Framework – MLVTG
Figure 2: Overview of the MLVTG Framework. The input video and query are first encoded by frozen feature encoders, then projected into a shared semantic space via FFN (Sec. 3.2.1). One branch computes saliency scores directly, while the other processes concatenated features through the MambaAligner (Sec. 3.2.2), where bidirectional 1D convolutions and SiLU-activated recurrent SSMs dynamically filter spatiotemporal features via selective memory. The aligned features undergo semantic purification through mamba-based LLMRefiner (Sec. 3.2.3), leveraging implicit multi-modal correlations to align modalities while preserving task-relevant semantics. Finally, task-specific heads (Sec. 3.2.4) classify the optimized features to produce grounding results.
As illustrated in Fig.2, MLVTG integrates the MambaAligner with LLMRefiner, optimizing a multi-modal feature alignment mechanism. Through a dual-branch decoupled task architecture, collaborative reasoning optimization is achieved for temporal localization and highlight detection.
3.2.1 Feature Extraction and Projection.
Given an input video with $L_{v}$ clips and a text query with $L_{q}$ tokens, we follow the encoder design of [23] to extract features, which are then projected into a shared $D$-dimensional space via separate Feed-Forward Networks. This yields video features $\mathbf{V}=\{\mathbf{v}_{i}\}_{i=1}^{L_{v}}\in\mathbb{R}^{L_{v}\times D}$ and text features $\mathbf{Q}=\{\mathbf{q}_{j}\}_{j=1}^{L_{q}}\in\mathbb{R}^{L_{q}\times D}$.
We propose a dual branch task decoupled architecture. In the one branch, an attentive pooling mechanism aggregates query tokens $\mathbf{Q}\in\mathbb{R}^{L_{q}\times D}$ into a sentence-level representation $\mathbf{S}\in\mathbb{R}^{1\times D}$ to compute a saliency score via similarity measurement:
$\mathbf{S}=\mathbf{M}\mathbf{Q},\quad\mathbf{M}=\text{Softmax}(\mathbf{W}\mathbf{Q})\in\mathbb{R}^{1\times L_{q}},$ (1)
where $\mathbf{W}\in\mathbb{R}^{1\times L_{q}}$ is a learnable embedding. In another branch, we incorporate positional embeddings $\mathbf{E}^{pos}$ and modality-type embeddings $\mathbf{E}^{type}$ into both video and query tokens to preserve spatial and modality distinctions:
$\tilde{\mathbf{V}}=\mathbf{V}+\mathbf{E}^{pos}_{V}+\mathbf{E}^{type}_{V},\quad\tilde{\mathbf{Q}}=\mathbf{Q}+\mathbf{E}^{pos}_{T}+\mathbf{E}^{type}_{T}.$ (2)
These representations are concatenated as $\mathbf{Z}=[\tilde{\mathbf{Q}};\tilde{\mathbf{V}}]\in\mathbb{R}^{L\times D}$, where $L=L_{q}+L_{v}$. Subsequently, $\mathbf{Z}$ passes through MambaAligner for robust video representations and multi-modal alignment.
3.2.2 MambaAligner.
To better capture video representations for multi-modal alignment, MambaAligner consists of $K$ stacked vision mamba blocks and takes the concatenated feature sequence $\mathbf{Z}=[\tilde{\mathbf{Q}};\tilde{\mathbf{V}}]$ as the input (see Fig. 2). The input is first normalized and projected into two branches:
$\hat{\mathbf{Z}}=\mathrm{Norm}(\mathbf{Z}),\quad\mathbf{x}=W_{x}\hat{\mathbf{Z}},\quad\mathbf{g}=W_{g}\hat{\mathbf{Z}},$ (3)
where $\mathrm{Norm}$ denotes layer normalization, and $W_{x}$, $W_{g}$ are learnable projections. Here, $\mathbf{x}$ is used for sequence modeling, while $\mathbf{g}$ serves as the gating signal.
A 1D convolution layer extracts local features from $\mathbf{x}$, generating forward and backward sequences $\mathbf{x}^{f}$ and $\mathbf{x}^{b}$, respectively. Each is passed through a SSM with parameters $(A,B,C)$:
$\mathbf{y}=\mathbf{x}*\mathbf{K},\quad\mathbf{K}=\begin{bmatrix}CA^{0}B,CA^{1}B,\cdots,CA^{L-1}B\end{bmatrix},$ (4)
where $\mathbf{K}$ is the impulse response kernel and * denotes convolution. This yields $\mathbf{y}^{f}$ and $\mathbf{y}^{b}$, capturing global context in both directions.
To fuse the outputs, a gating mechanism uses the control signal $\mathbf{g}$:
$\mathbf{y}_{\text{fused}}=\sigma(\mathbf{g})\odot\mathbf{y}^{f}+\big{(}1-\sigma(\mathbf{g})\big{)}\odot\mathbf{y}^{b},$ (5)
where $\sigma(\cdot)$ is the SiLU activation, and $\odot$ denotes element-wise multiplication. The final output is computed via residual connection:
$\mathbf{Z}_{\text{out}}=\mathbf{Z}+\mathbf{y}_{\text{fused}}.$
3.2.3 LLMRefiner.
To refine the semantics of latent representations and enhance multi-modal alignment, we propose the mamba-based LLMRefiner module. This module integrates two linear layers, denoted as $F_{L}^{1}$ and $F_{L}^{2}$, along with a pre-trained mamba-based large language model layer ($F_{LLM}$). Recent studies have highlighted the robust multi-modal reasoning abilities inherent to LLM, demonstrating their effectiveness in noise suppression and semantic refinement. Leveraging these inherent capabilities, our approach utilizes a specific frozen layer from pre-trained LLM. This mamba-based layer, $F{LLM}$, operates with frozen parameters to transfer textual priors effectively into the visual domain, thereby filtering out redundant or irrelevant video content.
Due to the dimension mismatch between the output of the MambaAligner($\mathbf{Z}{out}$), and the input dimension requirements of the $F_{LLM}$ layer, we integrate linear projection layers $F_{L}^{1}$ and $F_{L}^{2}$ before and after the $F_{LLM}$ block, respectively. These projection layers harmonize the dimensional compatibility, modifying the network output according to the following formulation:
$\mathbf{Z}_{refine}=F_{L}^{1}*F_{LLM}(\mathbf{Z}_{out})*F_{L}^{2}.$ (6)
In our implementation, the parameters of the pre-trained $F_{LLM}$ block remain entirely frozen to preserve their robust semantic representation capability. In contrast, the linear projection layers, $F_{L}^{1}$ and $F_{L}^{2}$, are fully trainable and optimized throughout the training process. This targeted training approach ensures continuous refinement of semantic alignment specifically tailored for downstream multi-modal alignment tasks.
3.2.4 Joint Optimization for TL & HD.
To jointly solve TL and HD, we design a dual-head module to produce task-specific outputs coupled with corresponding optimization objectives. Specifically, the TL Head takes refined feature $\mathbf{Z}_{refine}$ as input and integrates two parallel parts composed of 1D conv layers. One predicts the start and end timestamps ($[st,ed]$), while the other performs frame-level binary classification to discriminate foreground from background frames. In contrast, the HD Head receives video features $\mathbf{V}$ and sentence-level representations $\mathbf{S}$ as input and focuses on frame-level salient scores. The overall training objective aggregates task-specific losses with balancing coefficients, formulated as follows:
$\mathcal{L}_{overall}$ $=\mathcal{L}_{TL}+\mathcal{L}_{HD}$ (7)
$=\lambda_{f}\mathcal{L}_{f}+\lambda_{reg}\mathcal{L}_{reg}+\lambda_{inter}\mathcal{L}^{inter}+\lambda_{intra}\mathcal{L}^{intra},$
where $\mathcal{L}_{f}$ represents the classification loss, and $\mathcal{L}_{reg}$ denotes the regression loss, both dedicated to optimizing TL. Conversely, $\mathcal{L}^{inter}$ and $\mathcal{L}^{intra}$ are explicitly employed to enhance the accuracy and discriminative power of HD. Detailed formulations and implementation specifics of each loss term are elaborated in the supplementary materials.
Table 1: Performances on QVHighlights test and val splits. The optimal results are indicated in bold, and the suboptimal results are underlined.
[…]
Method TL HD TL HD
R1 mAP $\geq$ Very Good R1 mAP $\geq$ Very Good
@0.5 @0.7 @0.5 @0.75 Avg. mAP HIT@1 @0.5 @0.7 @0.5 @0.75 Avg. mAP HIT@1
MCN [14]ICCV’17 11.4 2.7 24.9 8.2 10.7 - - - - - - - - -
CAL [7]ICCV’22 25.5 11.5 23.4 7.7 9.9 - - - - - - - - -
XML+[19]ICCV’20 46.7 33.5 47.9 34.7 34.9 35.4 55.1 - - - - - - -
M-DETR [18]NeurIPS’21 52.9 33.0 54.8 29.4 30.7 35.7 55.6 53.9 34.8 - - 32.2 35.7 55.6
UMT [24]CVPR’22 56.2 41.2 53.4 37.0 36.1 38.2 60.0 60.3 44.3 - - 38.6 39.9 64.2
UniVTG [23]ICCV’23 58.9 40.9 57.6 35.6 35.5 38.2 61.0 59.7 - - - 36.1 38.8 61.8
MH-DETR [38]ICCV’22 60.0 42.4 60.7 38.1 38.3 38.2 60.5 60.8 44.9 60.7 39.6 39.2 38.7 61.7
QD-DETR [26]CVPR’23 62.4 45.0 62.5 39.9 39.9 38.9 62.4 62.7 46.7 62.2 41.8 41.2 39.1 63.0
TR-DETR [31]AAAI’23 64.6 48.9 63.9 43.7 42.6 39.9 63.4 - - - - - - -
TaskWeave [39]CVPR’24 - - - - - - 64.2 64.2 50.0 65.3 46.4 45.3 39.2 63.6
UVCOM [36]CVPR’24 63.5 47.4 63.3 42.6 43.1 39.7 64.2 65.1 51.8 - - 45.7 - -
[…]
4 Experiments
4.1 Datasets & Experimental Settings
4.1.1 Datasets.
We evaluate our method on three widely-used benchmarks: QVHighlights, Charades-STA, and TVSum. QVHighlights [18] is the first benchmark designed for joint evaluation of video temporal grounding and highlight detection. It consists of over 10,000 YouTube videos spanning diverse categories such as lifestyle vlogs, news, and documentaries. Our main experiments were conducted on this dataset. Charades-STA [9] augments the original Charades dataset by adding temporal annotations aligned with natural language. It comprises 9,848 indoor activity videos and 16,128 sentence-moment pairs. It is used to evaluate the model performance on TL. TVSum [30] includes 50 YouTube videos across 10 categories (e.g., DIY tutorials) and is used for HD evaluation.
4.1.2 Evaluation Metrics.
We adopt standard evaluation metrics specific to each dataset. For QVHighlights, we follow the official evaluation metric proposed in [19]. For TL, we report Recall@1 at IoU thresholds 0.5 and 0.7, mAP at 0.5 and 0.75, and the average mAP over 0.5–0.95 with a step size of 0.05. For HD, we use mAP and HIT@1, where a clip is counted as correct if its saliency is annotated as Very Good. For Charades-STA, we report Recall@1 at IoU 0.5, and 0.7, along with mIoU. For TVSum, we follow the evaluation metric in [30], reporting mAP and Top-5 mAP.
4.1.3 Implementation Details.
We concatenate CLIP (ViT-B/32)and SlowFast (R-50) features from pre-segmented video clips at fixed frame rates (0.5/1 FPS), while text features are extracted through the CLIP text encoder. We use $k{=}4$ vision mamba blocks with $d{=}1024$ hidden dimensions in the MambaAligner. The LLMRefiner operates on 2056 dimension features extracted from the 20th layer. We use a batch size of 32 and the model is trained for 200 epochs on a single RTX 4090 GPU. Details are provided in the supplementary materials.
4.2 Comparison with Start-of-the-Arts
4.2.1 Joint Temporal Localization and Highlight Detection.
Our MLVTG achieves state-of-the-art performance on TL and HD, as shown in Table 3.2.4. On the QVHighlights test set, it outperforms UVCOM by 0.8% in R1@0.7 and achieves the highest mAP@0.75 (43.9) and HD mAP (39.9). Since TR-DETR specifically designs a module utilizing HD to assist the TL on QVHighlights, it performs well. However, more generalized MLVTG outperforms it on the generic TL dataset Charades-STA, as shown in Table 2. On the val set, MLVTG achieves 65.1 R1@0.5 (+0.9% over UVCOM), 50.5 R1@0.7 (+8.6% over QD-DETR), and 46.8 mAP@0.75 (+0.9% over TaskWeave). For HD evaluation, it also reaches 65.2 HIT@1, surpassing UVCOM and QD-DETR by 2.5% and 3.5%.
Table 2: Performances on Charades-STA. The suboptimal results are underlined
Temporal Localization
R1
Method @0.5 @0.7 mIoU
CTRL [9] 23.6 8.8 -
MAN [43] 46.5 22.7 -
M-DETR [18] 53.6 31.3 -
VSLNet [44] 54.1 35.2 50.0
UniVTG [23] 58.0 35.6 50.1
QD-DETR [26] 55.5 34.1 -
TR-DETR [31] 57.6 33.5 -
TaskWeave [39] 56.5 33.6 -
UVCOM [36] 59.2 36.6 -
MLVTG 58.3 38.7 50.3
4.2.2 Temporal Localization.
In the TL task, we evaluate MLVTG on the Charades-STA, comparing its performance with advanced temporal localization methods (Table 2). The experiments measure R1@0.3, 0.5, 0.7 and mIoU. MLVTG achieves the best results on R1@0.7 and mIoU, surpassing UVCOM by 2.04% in R1@0.7, highlighting its robustness under strict metrics. Its R1@0.5 (58.3) closely matches UVCOM (59.2), while outperforming detector-based methods like TR-DETR (57.6). Notably, without relying on complex cascading detector designs, MLVTG achieves superior overall performance, indicating balanced localization accuracy and generalization capability.
4.2.3 Highlight Detection.
Table 3: Highlight Detection performance of Top-5 mAP on TVSum. “†” denotes the methods that utilize the audio modality. The suboptimal results are underlined.
Method Highlight Detection
VT VU GA MS PK PR FM BK BT DS Avg.
Trailer [33] 61.3 54.6 65.7 60.8 59.1 70.1 58.2 64.7 65.6 68.1 62.8
SL-Module [37] 86.5 68.7 74.9 86.2 79.0 63.2 58.9 72.6 78.9 64.0 73.3
PLD-VHD [35] 84.5 80.9 70.3 72.5 76.4 87.2 71.9 74.0 74.4 79.1 77.1
UniVTG [23] 83.9 85.1 89.0 80.1 84.6 81.4 70.9 91.7 73.5 69.3 81.0
MINI-Net [15] 80.6 68.3 78.2 81.8 78.1 65.8 57.8 75.0 80.2 65.5 73.2
TCG [40] 85.0 71.4 81.9 78.6 80.2 75.5 71.6 77.3 78.6 68.1 76.8
Joint-VA [2] 83.7 57.3 78.5 86.1 80.1 69.2 70.0 73.0 97.4 67.5 76.3
CO-AV [21] 90.8 72.8 84.6 85.0 78.3 78.0 72.8 77.1 89.5 72.3 80.1
MLVTG 83.6 85.6 74.7 75.9 82.4 87.7 64.6 91.6 80.9 73.7 80.1
On TVSum (Table 3), MLVTG achieves an average F1-score of 80.1, on par with the state-of-the-art CO-AV [21], and excels in fine-grained scenarios. It notably outperforms CO-AV in user-generated content (VU: 85.6, +12.8%) and ads (PR: 87.7, +9.7%). UniVTG uses multiple datasets for training. Even without this, MLVTG remains competitive. MLVTG demonstrates superior generalization capabilities over complex methods like Joint-VA without introducing audio modality, highlighting the advantages of our method.
5 Ablation Study
In this section, we conduct ablation studies on three aspects and detailed experimental results are provided in the supplementary materials.
As shown in Table 4, MambaAligner alone significantly boosts TL results—on QVHighlights, R1@0.7 improves by +5.4%, mIoU by +4.4%, and mAP@Avg by +5.3%; on Charades-STA, gains are +3.1%, +0.8%, and +1.4%. LLMRefiner contributes moderate improvements (e.g., +1.6% R1@0.7 on QVHighlights), reflecting its semantic refinement effect. Combining both yields the best TL results, with total gains of +7.0% R1@0.7, +5.2% mIoU, and +5.6% mAP@Avg. On HD, MambaAligner provides +0.8% mAP and +4.4% Avg, LLMRefiner adds minor benefits, and the full model achieves the highest HD performance (+1.2% mAP, +7.4% Avg), confirming their complementarity.
Table 4: Ablation Study on Three Benchmarks. We select the most representative indicators as validation metrics for each task.
Mamba Aligner LLM Refiner Temporal Localization Highlight Detection
QVHighlights Charades-STA QVHighlights TVSum
R1 R5 mIoU mAP R1 mIoU mAP $\geq$ Very Good Avg.
@0.7 @0.7 @0.75 Avg. @0.7 @0.75 Avg HIT@1 mAP
[…]
Figure 3: Effectiveness of LLM Pretrained Parameters.” means no LLM layer. “Rand” means randomly initialized.
LLM Param Frozen R1 R5 mAP
@0.7 @0.7 @0.75 @Avg.
43.5 59.4 39.0 37.7
Rand No 43.6 57.4 37.0 36.7
Rand Yes 43.3 57.9 38.1 37.1
Yes No 42.8 57.2 36.9 36.4
Yes Yes 45.1 61.4 39.3 38.4
Figure 4: Effectiveness of LLM Different Layer.
To validate the effectiveness of frozen pre-trained LLM parameters, we conduct experiments on the baseline with different initialization and freezing strategies. All models adopt a single-layer Transformer for fair comparison (Table 5). The baseline achieves 38.4 mAP@Avg. Adding a trainable random initialized layer reduces performance by 2.7% (row 2 vs. 1), showing that structural imitation introduces noise. Freezing random initialized layers yields 37.1 (–3.4%, row 3 vs. 5), highlighting the need for pre-trained LLM parameters. Full fine-tuning leads to 36.4 (–5.2%, row 4 vs. 5), confirming that updates disrupt alignment. In contrast, frozen LLMs achieve 45.1 R1@0.7, 38.4 mAP@Avg, and 61.4 R5@0.7, demonstrating better temporal grounding.
Layer-wise analysis reveals that the multi-modal alignment ability of LLMs is highly dependent on layer selection. As shown in Fig 4, using the 20th layer’s frozen parameters yields the best performance. In contrast, higher layer (e.g., 25th) lead to significant degradation, suggesting that excessive abstraction harms visual grounding. Overall, performance shows a parabolic trend across layers, with the range 15–22th forming a “semantic sweet spot” that avoids both syntactic noise from lower layers and semantic oversimplification from higher ones.
6 Visualization
Figure 5: Feature Alignment Visualization. Darker colors indicate lower similarity.
Figure 5 illustrates the cosine similarity heatmap between query and video features. Initially, the alignment between these features is poor, as evidenced by the scattered distribution of similarity scores. Upon introducing MambaAligner, a substantial improvement in the alignment quality between video and textual features is observed. Nevertheless, the presence of spurious regions exhibiting high similarity, particularly in the early video clip due to background distractors, reveals limitations in alignment precision. Subsequent refinement by LLMRefiner markedly enhances alignment accuracy by effectively suppressing noise and concentrating high-similarity regions within the ground-truth interval (63–132 seconds). The visualization explicitly demonstrates the incremental benefit provided by MambaAligner for enhancing alignment and the critical refinement contributed by LLMRefiner for precise localization. Furthermore, Fig. 6 further compares our method MLVTG with the baseline on QVHighlights.
Figure 6: Visualization of joint temporal localization and highlight detection on QVHighlights.
7 Conclusion
This paper proposes MLVTG, a novel framework for video temporal grounding that combines Mamba and frozen LLM layers for effective multi-modal alignment. Experiments on three benchmarks show that MLVTG outperforms state-of-the-art methods in both temporal localization and highlight detection. Ablation studies further confirm the effectiveness of each component, demonstrating the robustness and superiority of MLVTG. Future work will explore extending MLVTG to incorporate audio modalities for more comprehensive video temporal grounding.